OpenStudy (anonymous):

4^2 • 4^3 Kerry is trying to solve the above expression. She takes the following steps: Kerry notices the bases of each power are the same, so she writes 4 down as the base of her answer. Kerry sees that this is a multiplication problem, so she multiplies the exponents together to get 6 as the exponent of her answer. Kerry arrives at the answer 4^6. Which step(s) did she make a mistake in and what was the mistake?

OpenStudy (saifoo.khan):

When bases are same and there's a multiply sign in b/w you add the exponents.

OpenStudy (shubhamsrg):

mistake in the very first step -> kerry attempts to solve this problem! :|

OpenStudy (anonymous):

\[x^a \times x^b = x^{a+b}\]

OpenStudy (anonymous):

She did mistake in second step by multiplying the two exponents...

OpenStudy (ganpat):

it should be...\[4^{2} * 4^{3} = 4 ^{(2 +3)} = 4^{5} = 1024\]